A car traveled for 1/2 hour at r mi/h, then increased the speed by 10 mi/h and traveled for 1 1/2 hours more. How far did the car go? (Express the answer in terms of r)
Answer by stanbon(48550)

A car traveled for 1/2 hour at r mi/h, then increased the speed by 10 mi/h and traveled for 1 1/2 hours more. How far did the car go? (Express the answer in terms of r)
----------
1st half-hour DATA:
Rate = r mph ; Time = 1/2 hr ; distance = rt = (1/2)r miles
-------------
Next 1 1/2 hr DATA:
Rate = (r+10) mph ; Time = (3/2) hr ; Distance = rt = (r+10)*(3/2)
-------------------
Total Distance
= (1/2)r + (3/2)r + 15
= (2r+15) miles
=========================
Cheers,
Stan H.